OpenStudy (anonymous):

The velocity of a particle moving along the x-axis is v(t) = t2 – 2t, with t measured in minutes and v(t) measured in feet per minute. To the nearest foot find the total distance travelled by the particle from t = 0 to t = 3 minutes.

OpenStudy (anonymous):

@zepdrix pretty sure I just integrate, just wanted to confirm

zepdrix (zepdrix):

Ummm

OpenStudy (anonymous):

https://gyazo.com/1ad2f2f534c3e4c134c1e809420a3e7a

OpenStudy (welshfella):

yes integrate between t = 0 and t = 3

zepdrix (zepdrix):

I mean, yes, integrating will get you your position function, good. The issue is, if the particle moved backwards and then also forwards during those first 3 minutes, then we might have an issue. The s(t) is going to tell us about displacement, not total distance. So maybe we need to look at critical points first.

zepdrix (zepdrix):

Maybe I'm mistaken hehe, rusty

OpenStudy (anonymous):

It says move along so I doubt it means back and forth.

OpenStudy (anonymous):

Anyways I got 0.

OpenStudy (welshfella):

yes it does move in 2 directions 0 is the displacement no the distance travelled

zepdrix (zepdrix):

Setting v(t)=0 gives us:\[\large\rm 0=t^2-2t\qquad\to\qquad t=0,~t=2\] These are the times when the velocity changes from positive to negative. If you do a sign chart or something similar, you see that the particle was moving backwards from t=0 to t=2. So maybe what you need to do is break up the integral.

OpenStudy (welshfella):

yes

zepdrix (zepdrix):

\[\large\rm total~distance=\left|\int\limits_0^2 t^2-2t~dt\right|+\int_2^3 t^2-2t~dt\] Somethingggg like that.

OpenStudy (welshfella):

yes

zepdrix (zepdrix):

Maybe that's more complicated :) Just integrate,\[\large\rm s(t)=\frac{1}{3}t^3-t^2+c\] The particle is moving backwards from t=0 to t=2, so calculate the distance it covered in that interval,\[\large\rm s(2)-s(0)\]And then calculate the distance it traveled while moving forward,\[\large\rm s(3)-s(2)\]and add those quantities together. Same thing, just without the ugly integrals I guess.

OpenStudy (anonymous):

Nah I like the first one better, actually how I remember doing it. I got 2.667

OpenStudy (welshfella):

the graph looks something like below so you need to calculate the 2 areas